How much should a couple invest at the time their son is born to provide him with $50,000 for his college tuition. If the interest is 3.5% compounded quarterly(Assume he is 18 when he goes to college)
Answer by stanbon(75887)

How much should a couple invest at the time their son is born to provide him with $50,000 for his college tuition. If the interest is 3.5% compounded quarterly(Assume he is 18 when he goes to college)
----------------------
A(t) = P(1+(r/n))^(n*t)
----------------------------------
50,000 = P(1+0.035/4)^(4*18)
-----
50,000 = P(1.00875)^(72)
----
50,000 = P(1.8725)
----
Ans: P = $26,702.27
--------------------
Cheers,
Stan H.
---------------